OpenStudy (anonymous):

Chico walks 3/5 mile to the library. Nikki walks 1/4 the distance that Chico walks to the library. How far does Nikki walk?

OpenStudy (anonymous):

Not that far actually

OpenStudy (boldjon):

lol that answer though XD

OpenStudy (anonymous):

wat do u mean ?

OpenStudy (anonymous):

Just kidding. @dulceceron You simply have to multiply both fractions to get your answer.\[\frac{ 3 }{ 5 } \times \frac{ 1 }{ 4 } = ~?\]

OpenStudy (anonymous):

is it 3/20

OpenStudy (boldjon):

U GOT IT

OpenStudy (boldjon):

awesome work

OpenStudy (anonymous):

thank you @boldjon